Rachel travelled by taxi from her home to the hospital.
The journey from her home to the hospital was 18 km.
Here is the formula for calculating a taxi fare.
F= 320 + 30d
F is the fare in pence
d is the distance travelled in kilometres
The taxi driver charged Rachel £10
Was she overcharged?

Answer by fractalier(6550)

From the formula
F = 320 + 30d
we plug in the number of kilometers...
F = 320 + 30(18) = 320 + 540 = 860 pence
If 10 pounds is 1000 pence, she was overcharged 140 pence...